Find the annual percentage rate for an investment that earns 6% per year, compounded quarterly
Answer by stanbon(75887)

Find the annual percentage rate for an investment that earns 6% per year, compounded quarterly
-----
P(1+(0.06/4)^4
---
= P(1.015)^4
= P(1.0614)
= P + 0.0614P
----
Annual percentage rate = 6.14%
===================================
Cheers,
Stan H.